Not sure how to go about figuring out the energy conversion for this

1) If you were to send a ball up a ramp at twice the speed of a previous trial, how much higher could it go up the ramp?
Answer #1

let v is the initial speed and h is height raised by the ball.

Apply conservation of energy

final gravitational potential energy = initial kinetic energy

m*g*h = (1/2)*m*v^2

h = v^2/(2*g)

let h' is the height raised when speed is doubled

h' = (2*v)^2/(2*g)

= 4*(v^2/(2*g))

= 4*h <<<<<------Answer

so, the object can go 4 times of the initial height.